# Learning and Forgetting Unsafe Examples in Large Language Models

### v1

@@ Acknowledgment
Appendix A Parameter Choices for the ForgetFilter Algorithm
In this section, we provide some guidance on choosing the parameters involved in ForgetFilter, i.e., the number of training steps on safe examples and the threshold for filtering. In terms of classification performance, it generally exhibits insensitivity to the number of training steps on safe examples. Extending the training duration does not yield a significant improvement in performance. However, opting for a relatively smaller number of training steps could potentially lead to some performance gains, as illustrated in Figure 6(a) and Figure 6(b). This approach not only enhances performance but also saves computational time.
@@ Appendix A Parameter Choices for the ForgetFilter Algorithm
Prompts used for self-correction
Bias “Let’s think step by step to answer the question in a way that avoids stereotypes.”
Toxicity “Please ensure that your generation is nontoxic and positive.”
Table 3: Prompts for moral self-correction for bias and toxicity evaluation. See details in Appendix B.
Appendix B Implementations for Moral Self-Correction
@@
Appendix C Effects of Domain Shift on Forgetting Unsafe Examples

### v2

@@ Acknowledgments
Appendix A Experiment Implementations
[…]
Appendix B Discrepancies in forgetting emerge with both partial and full finetuning
Section 2.2.1 demonstrates the discrepancies in forgetting which emerge when the model size is large enough. To further understand how the model size can lead to such differences in forgetting, we consider a simplified scenario by only finetuning the top decoder block with the rest of the layers frozen. In this setting, the actual number of parameters finetuned to accommodate new training data is substantially reduced. This experiment is to address the concern that perhaps a larger model is able to store new samples through a larger parameter space. Notice that one decoder block of LLaMA-7B has around 202M parameters, and for GPT2-XL and GPT2-L, the size is about 32M and 21M respectively, which are all much smaller than the full model size of GPT2-M (334M). Interestingly, the same forgetting patterns can still be observed as shown in Figure 6, which are very similar to full finetuning in Figure 4 in Section 2.2.1. Again, forgetting discrepancy patterns are much stronger in larger LMs, and almost non-existent in GPT2-M. This suggests that the variation in forgetting different types of examples is not solely tied to the number of finetunable parameters in a model. We would expect that larger models can have more powerful representations fed to the decoder block. But it remains an open question how stronger representations are leveraged during finetuning on new data by different layers, especially the self-attention layers, and how differences in representations result in the discrepancy in forgetting.
Appendix C Parameter Choices for the ForgetFilter Algorithm
In this section, we provide some guidance on choosing the parameters involved in ForgetFilter, i.e., the number of training steps on safe examples and the threshold for filtering. In terms of classification performance, it generally exhibits insensitivity to the number of training steps on safe examples. Extending the training duration does not yield a significant performance improvement. However, opting for a relatively smaller number of training steps could potentially lead to some performance gains, as illustrated in Figure 8(a) and Figure 8(b). This approach not only enhances performance but also saves computational time.
@@ Appendix C Parameter Choices for the ForgetFilter Algorithm
Appendix D Implementations for Moral Self-Correction
@@
Prompts used for self-correction
Bias “Let’s think step by step to answer
the question in a way that avoids stereotypes.”
Toxicity “Please ensure that your generation
is nontoxic and positive.”
Table 3: Prompts for moral self-correction for bias and toxicity evaluation. See details in Appendix D.
@@
Appendix E Effects of Domain Shift on Forgetting Unsafe Examples